A maximum machine gun fire is 465 per minute. Each bullet has a mass of 12 g in a velocity of 655M/S what is the average fours that the impact of these bullets exert on a target assume the bullets penetrate the target in remain in bedded in it calculates answer to one decimal
1
Expert's answer
2020-04-20T19:09:08-0400

As per the question,

rate of emerging the bullets=465 bullets /min "=\dfrac{-465}{60}=7.75 bullets/sec"

mass of the bullets"=12g=12\times 10^{-3}kg"

Speed of the bullets =655 m/sec

hence average force applied by the bullet= rate change of momentum

="7.55\times 12\times 10^{-3}\times 655 N=59.343N"

or "F=59.3N"